OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7.

OpenStudy (anonymous):

y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14 answer choices

OpenStudy (anonymous):

First, determine the slope of the new line. Remember that lines that are perpendicular have slopes that are negative reciprocals of each other. So, if the new line is perpendicular to a line with a slope of 1/3, what is the slope of the new line?

OpenStudy (anonymous):

-3/1?

OpenStudy (anonymous):

Right, which is just -3 right? So now, the equation we'll be using here is: \[y-y_1=m(x-x_1)\]Where m is the slope of the line, and x_1,y_1 are the coordinates of any point on the line. We have the slope and a point on the line, so you can plug in and solve.

OpenStudy (anonymous):

y- -2 = -3(x - 4) y - -2 = -3x - 12 + 2 +-2 y = -3x + 10

OpenStudy (anonymous):

I mean y = -3x - 10

OpenStudy (anonymous):

y--2=-3x-12 y+2=-3x-12 y=-3x-14

OpenStudy (anonymous):

thank you! can you help me with another question ?